# Visual Diffusion Models are Geometric Solvers

### v1

@@
Abstract.
In this paper we show that visual diffusion models can serve as effective geometric solvers: they can directly reason about geometric problems by working in pixel space. We first demonstrate this on the Inscribed Square Problem, a long-standing problem in geometry that asks whether every Jordan curve contains four points forming a square. We then extend the approach to two other well-known hard geometric problems: the Steiner Tree Problem and the Simple Polygon Problem. Our method treats each problem instance as an image and trains a standard visual diffusion model that transforms Gaussian noise into an image representing a valid approximate solution that closely matches the exact one. The model learns to transform noisy geometric structures into correct configurations, effectively recasting geometric reasoning as image generation. Unlike prior work that necessitates specialized architectures and domain-specific adaptations when applying diffusion to parametric geometric representations, we employ a standard visual diffusion model that operates on the visual representation of the problem. This simplicity highlights a surprising bridge between generative modeling and geometric problem solving. Beyond the specific problems studied here, our results point toward a broader paradigm: operating in image space provides a general and practical framework for approximating notoriously hard problems, and opens the door to tackling a far wider class of challenging geometric tasks.
**footnotetext: Equal contribution.
1. introduction
@@ Abstract.
Diffusion models have been used in various contexts to tackle optimization and reasoning problems, including combinatorial tasks such as the traveling salesman problem  (Sun & Yang, 2023; Li et al., 2023; Sanokowski et al., 2025). These approaches typically formulate the problem in symbolic or graph-based representations, leveraging the probabilistic nature of diffusion to search solution spaces. In contrast, our method operates purely in the visual domain. By representing geometric problems as images and reasoning directly in pixel space, we exploit the intrinsic strength of diffusion models in handling multimodal distributions and ambiguous solutions. This visual formulation makes our approach fundamentally distinct from prior problem-solving applications of diffusion.
Figure 1. We introduce a visual diffusion approach to solving hard geometric problems directly in pixel space. Shown here on the Inscribed Square Problem, where we task the model with finding a square such that all of its four vertices lie on a given curve. Our method uncovers diverse approximate solutions, corresponding to different random seeds.
To ground our approach, we begin with the Inscribed Square Problem, a long-standing problem that asks whether every simple closed curve in the plane admits an inscribed square. The problem is still unsolved in the general case. Furthermore, a given curve may admit multiple and often very different inscribed squares, and enumerating them is non-trivial even in restricted settings  (van Heijst, 2014). This multiplicity naturally forms a distribution, which makes the problem especially well suited to diffusion models. Our method addresses it in an unexpected way, operating directly in image space on a visual representation of the geometric challenge. By starting from different random seeds, the model can uncover diverse valid squares, each corresponding to a distinct solution of the problem. Figure 1 illustrates the setting and highlights both the complexity and the variability of possible solutions.
We next illustrate the competence of our diffusion-in-image-space approach on two additional hard geometric problems. The first is the Steiner Tree Problem, which asks for the shortest possible network connecting a given set of points. Its solution may introduce auxiliary nodes, known as Steiner points, and finding the optimal configuration is NP-hard  (Garey et al., 1977). The second is the problem of connecting a set of points into a simple polygon of maximum area, which was featured in the CG:SHOP global optimization challenge of 2019  (Demaine et al., 2022). This task is known for its combinatorial complexity and strict geometric constraints, and is also NP-hard. As with the inscribed square problem, our method addresses these problems directly in image space, operating in the pixel domain. While discretization at finite resolution imposes limitations, it nonetheless enables valid approximations to problems that are otherwise extremely hard, with solutions that can be naturally refined. We evaluate this aspect rigorously in the paper.
@@ 2. Related Work
Several works have attempted to use diffusion models in order to learn to solve problems for which no known efficient algorithm exists, mostly of combinatorial nature. Most existing diffusion works operate directly on the parameter space of the problem. DIFUSCO  (Sun & Yang, 2023), a graph-based diffusion framework, casts a broad family of NP-complete problems into $\{0,1\}^{N}$ indicator vectors and learns a denoiser over graphs. They compare continuous (Gaussian) and discrete (Bernoulli) noise processes, and show strong results on traveling salesman problem (TSP) and maximum independent set (MIS). Complementing the supervised approach, the T2T  (Li et al., 2023) line of work learns a distribution of high-quality solutions during training and then performs gradient-guided optimization at test time by iteratively noising the current solution and denoising while guiding towards lower energy solutions of some relaxed objective, achieving competitive quality–efficiency trade-offs on TSP and MIS. Fast T2T  (Li et al., 2024) significantly speeds this up via an optimization-consistency objective, matching or surpassing multi-step diffusion solvers performance with single-step generation plus a single gradient step.
Closer to our approach, some methods solve constrained problems in pixel-space representation. (Graikos et al., 2023) train an unconditional diffusion model on pixel-space representations of TSP instances. They then solve new instances with stochastic optimization using a differential renderer, utilizing the prior of the learned model. Differently from us, they optimize a parametric representation of the solution to a given instance, while we generate solutions with DDIM sampling of a conditional model. (Wewer et al., 2025) train a noise-prediction UNet in pixel space on a visual representation of Sudoku, which is another NP-hard combinatorial problem. They depart from fully-parallel image diffusion by (i) assigning individual noise levels to patches, and (ii) sampling patches in a learned order or a hand-crafted order. They demonstrate that sampling order matters, and can substantially outperform a conditional DDPM baseline that operates fully in parallel.
@@
Figure 2. Example of a curve (black) with three inscribed squares. Note that the inscribed squares are defined only by having all four vertices on the curve: they need not be fully contained within the curve, and they may overlap with each other.
@@
Figure 3. Inscribed square $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed (inscribed square). Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; penultimate: t=0t{=}0). For $t\neq 0$ we render only the filled mask; at t=0t{=}0 we also draw square edges and the minimum-area bounding box. The rightmost column (“snapped”) shows the rigidly snapped version of the t=0t{=}0 prediction on the curve, with an arrow from the original centroid to the snapped centroid.
3. Inscribed Square Problem
@@
The problem has since been resolved in several restricted settings. Some works show the conjecture holds for convex and piecewise analytic curves (Emch, 1913, 1915, 1916) and later results prove it for $C^{1}$-smooth curves and for curves of finite total curvature or generic $C^{1}$ curves (Stromquist, 1989; Cantarella et al., 2021). More recent work demonstrates validity under additional low-regularity assumptions (Tao, 2017), yet it remains open for the general Jordan curve case.
@@
Algorithmic approaches for finding inscribed squares exist primarily for discrete or polygonal cases. For convex polygons, efficient procedures have been developed to detect or enumerate inscribed squares in subquadratic time (Chazelle, 1983; Sharir & Toledo, 1994). Discrete relaxations on grid polygons have likewise been studied and verified computationally for bounded cases (Pettersson et al., 2014). All these methods represent the curve symbolically or combinatorially, relying on exact geometric descriptions. In contrast, our work formulates the problem directly in pixel space.
@@
We address the inscribed square problem by formulating it entirely in image space and training a conditional diffusion model to recover inscribed squares from noisy inputs. To this end, we generated a dataset of synthetic samples. Each sample consists of a smooth Jordan curve $C$ together with one to five inscribed squares. The curves were constructed procedurally such that they pass through the square vertices while avoiding self-intersections, and both curves and squares were rasterized into $128\times 128$ binary images. Each training example pairs a curve image with one inscribed square. Full details of the curve-generation process are provided in the supplementary material, under Section B.1.
Our model follows the standard diffusion framework with a U-Net (Ronneberger et al., 2015) backbone with self-attention layers (Vaswani et al., 2017), similar to those employed in text-to-image diffusion models such as Stable Diffusion (Rombach et al., 2022). The conditioning signal is the clean binary image of the curve, while the ground truth $x_{0}$ is the clean image of the square, both represented in two-dimensional pixel space. The conditioning image is concatenated as an additional channel to the noisy input $x_{t}$ at each timestep, effectively treating the curve as a non-noisy color channel. We train with 100 denoising steps using a standard noise schedule and mean-squared error objective, enabling the model to transform noisy square images into valid inscribed squares consistent with the given curve.
@@
Figure 4. Solutions produced by our model. Each Jordan curve (black) is accompanied by predicted inscribed squares (colored).
(1) $\mathcal{A}(S,C)\;=\;-\frac{1}{4}\sum_{p\in V(S)}\operatorname{dist}(p,C),$
@@
We then apply a small rigid transformation $(R_{\theta},t)$ to $\widehat{S}$ and select the configuration that maximizes this alignment score. In practice, we approximate this optimization with a coarse grid search over small rotations and translations (see Section B.2 for details). This procedure nudges the predicted square so that its corners sit more closely on $C$, as visualized in the rightmost column of Fig. 3.
@@
(2) $\mathcal{Q}(S)\;=\;\frac{\operatorname{area}(S)}{w\cdot h}\;\cdot\;\exp\!\left(-\,2\,\Bigl|\tfrac{\max(w,h)}{\min(w,h)}-1\Bigr|\right).$
Table 1. Evaluation Results. We report alignment $\mathcal{A}(S,C)$ and squareness $\mathcal{Q}$ under three conditions: before snapping, after snapping, and ground truth (GT).
@@
4. Steiner Tree Problem
The second problem we cover in our case study is that of the Steiner Tree Problem. The Steiner tree problem asks, given a set of terminal points, to find a network of minimum total length that connects all terminals, where the construction is allowed to introduce additional points (Steiner points) to reduce length. In the Euclidean variant, which we focus on, Steiner points may be placed anywhere in the plane. The Steiner formulation is central to many applications where minimizing connection cost is critical. Some typical use cases for it include telecommunication  (Voss, 2006), PCB routing  (Chu & Wong, 2008), as well as infrastructure layout (roads, pipelines)  (Schwartz & Stückelberger, 2008; Cui et al., 2021).
Figure 5. Example of a Steiner Minimal Tree. Left: The input terminal nodes, colored in red. Right: The Steiner Minimal Tree for this instance, where auxiliary Steiner points are colored in dark gray.
@@ 4. Steiner Tree Problem
Formally, given a finite set of terminals $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{2}$, the Euclidean Steiner Tree (EST) problem asks for a straight-line embedded tree $T=(V,E)$ with $P\subseteq V$ that minimizes total Euclidean length $L(T)=\sum_{uv\in E}\|u-v\|_{2}$. Vertices in $V\setminus P$ are Steiner points. An optimal solution is called a Steiner minimal tree (SMT) and its length is denoted $L^{\star}(P)$ (Gilbert & Pollak, 1968). Figure  5 illustrates an instance of the problem and its optimal solution.
@@
For general $n$ (the number of initial given points), the EST problem is NP-hard already in the plane (Garey et al., 1977). However, there is a polynomial-time approximation scheme (PTAS) for Euclidean Steiner tree in fixed dimensions: for any fixed $\varepsilon>0$ one can compute a $(1+\varepsilon)$-approximate tree in time polynomial in $n$ (for fixed $\varepsilon$ and dimension).
@@
SMTs in the plane are highly structured: (i) no two edges cross; (ii) each Steiner point has degree exactly $3$ and the three incident edges meet at $120^{\circ}$; (iii) all angles in the tree are at least $120^{\circ}$ (at terminals that have degree >1>1); (iv) the number of Steiner points is at most n2n-2; and (v) all Steiner points lie in the convex hull of $P$ (Hwang & Richards, 1992; Brazil et al., 2004). These properties provide strong geometric constraints that are exploited by both exact and approximate methods, as discussed next.
@@
Exact algorithms typically rely on generating locally optimal full Steiner trees and then selecting a minimum-length subset, with implementations such as GeoSteiner solving large 2D instances (Warme et al., 2000; Juhl et al., 2018). On the approximation side, PTASes based on dissection or guillotine methods provide near-optimal guarantees in fixed dimensions (Arora, 1998; Mitchell, 1999).
@@
Figure 6. Steiner tree $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed. Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; rightmost: t=0t{=}0). Input points are overlaid in red.
While classical exact/approximate algorithms dominate practice, there have been several attempts to use learning-based solvers for the problem. For the Euclidean case, Wang et al. propose Deep-Steiner, which casts SMT construction as a sequential decision process: it discretizes the continuous search space, prunes candidates via KNN/MST neighborhoods, and then uses an attention-based policy trained with REINFORCE to add Steiner points iteratively (Wang et al., 2022). For non-Euclidean variants (rectilinear SMT and graph STP), several learning-based methods have been explored, including RL and GNN-driven solvers and mixed neural–algorithmic pipelines (Liu et al., 2021, 2024; Kahng et al., 2023; Ahmed et al., 2021; Du et al., 2021; Park et al., 2025).
@@
We generate a synthetic dataset by sampling random points (between 10 and 20 for each instance), and finding their SMT using the GeoSteiner solver (Juhl et al., 2018). Each solution is then rasterized into a grayscale image with different values for edges, nodes and background. The full details for data generation are found in Section B.3. We employ the same U-Net backbone as in the inscribed square experiment. The conditioning input consists of the rasterized terminal points, concatenated as an additional channel to the noisy input $x_{t}$ at each denoising step.
Figure 7. Optimal solutions (left) vs. our model’s solutions (middle) and the difference between them (right). Input points are overlaid over both optimal and produced solutions as red circles.
@@
Table 2. Steiner Tree Evaluation Results. Comparison of our method, MST, and random solutions. Reported are valid tree rates and mean Euclidean length ratios ($\pm$ std) relative to the optimal solution across input point ranges.
#Input Points Valid Rate Ours Ratio Mean $\pm$ Std MST Ratio Mean $\pm$ Std Random Ratio Mean $\pm$ Std
10–20 0.996 1.0008 $\pm$ 0.0005 1.0363 $\pm$ 0.0124 1.8344 $\pm$ 0.2363
21–30 0.986 1.0018 $\pm$ 0.0011 1.0416 $\pm$ 0.0095 1.9044 $\pm$ 0.1827
31–40 0.834 1.0044 $\pm$ 0.0035 1.0470 $\pm$ 0.0079 1.8981 $\pm$ 0.1656
41–50 0.334 1.0092 $\pm$ 0.0055 1.0522 $\pm$ 0.0072 1.8605 $\pm$ 0.1425
Figure 8. Qualitative examples of maximum area polygons (left) vs. polygons produced by our model (middle) and the difference between them (right). Areas depicted in red and blue in the difference map correspond to regions that are exclusive to the optimal solution and our solution, respectively. It can be noticed that even in cases where there is a disparity between the optimal solution and the one produced by the model, the area difference between the exclusive regions tends to be small, amounting to a solution of similar quality. Input points are overlaid over both optimal and produced solutions as red circles.
5. Maximum Area Polygon Problem

### v2

@@
Abstract
In this paper we show that visual diffusion models can serve as effective geometric solvers: they can directly reason about geometric problems by working in pixel space. We first demonstrate this on the Inscribed Square Problem, a long-standing problem in geometry that asks whether every Jordan curve contains four points forming a square. We then extend the approach to two other well-known hard geometric problems: the Steiner Tree Problem and the Simple Polygon Problem. Our method treats each problem instance as an image and trains a standard visual diffusion model that transforms Gaussian noise into an image representing a valid approximate solution that closely matches the exact one. The model learns to transform noisy geometric structures into correct configurations, effectively recasting geometric reasoning as image generation. Unlike prior work that necessitates specialized architectures and domain-specific adaptations when applying diffusion to parametric geometric representations, we employ a standard visual diffusion model that operates on the visual representation of the problem. This simplicity highlights a surprising bridge between generative modeling and geometric problem solving. Beyond the specific problems studied here, our results point toward a broader paradigm: operating in image space provides a general and practical framework for approximating notoriously hard problems, and opens the door to tackling a far wider class of challenging geometric tasks. Code is available at: https://kariander1.github.io/visual-geo-solver/.
Figure 1: We introduce a visual diffusion approach to solving hard geometric problems directly in pixel space. Shown here on the Inscribed Square Problem, where we task the model with finding a square such that all of its four vertices lie on a given curve. Our method uncovers diverse approximate solutions, corresponding to different random seeds.
1 introduction
@@ Abstract
Diffusion models have been used in various contexts to tackle optimization and reasoning problems, including combinatorial tasks such as the traveling salesman problem  [42, 27, 38]. These approaches typically formulate the problem in symbolic or graph-based representations, leveraging the probabilistic nature of diffusion to search solution spaces. In contrast, our method operates purely in the visual domain. By representing geometric problems as images and reasoning directly in pixel space, we exploit the intrinsic strength of diffusion models in handling multimodal distributions and ambiguous solutions. This visual formulation makes our approach fundamentally distinct from prior problem-solving applications of diffusion.
To ground our approach, we begin with the Inscribed Square Problem, a long-standing problem that asks whether every simple closed curve in the plane admits an inscribed square. The problem is still unsolved in the general case. Furthermore, a given curve may admit multiple and often very different inscribed squares, and enumerating them is non-trivial even in restricted settings  [44]. This multiplicity naturally forms a distribution, which makes the problem especially well suited to diffusion models. Our method addresses it in an unexpected way, operating directly in image space on a visual representation of the geometric challenge. By starting from different random seeds, the model can uncover diverse valid squares, each corresponding to a distinct solution of the problem. Figure 1 illustrates the setting and highlights both the complexity and the variability of possible solutions.
We next illustrate the competence of our diffusion-in-image-space approach on two additional hard geometric problems. The first is the Steiner Tree Problem, which asks for the shortest possible network connecting a given set of points. Its solution may introduce auxiliary nodes, known as Steiner points, and finding the optimal configuration is NP-hard  [17]. The second is the problem of connecting a set of points into a simple polygon of maximum area, which was featured in the CG:SHOP global optimization challenge of 2019  [9]. This task is known for its combinatorial complexity and strict geometric constraints, and is also NP-hard. As with the inscribed square problem, our method addresses these problems directly in image space, operating in the pixel domain. While discretization at finite resolution imposes limitations, it nonetheless enables valid approximations to problems that are otherwise extremely hard, with solutions that can be naturally refined. We evaluate this aspect rigorously in the paper.
@@ 2 Related Work
Several works have attempted to use diffusion models in order to learn to solve problems for which no known efficient algorithm exists, mostly of combinatorial nature. Most existing diffusion works operate directly on the parameter space of the problem. DIFUSCO  [42], a graph-based diffusion framework, casts a broad family of NP-complete problems into $\{0,1\}^{N}$ indicator vectors and learns a denoiser over graphs. They compare continuous (Gaussian) and discrete (Bernoulli) noise processes, and show strong results on traveling salesman problem (TSP) and maximum independent set (MIS). Complementing the supervised approach, the T2T  [27] line of work learns a distribution of high-quality solutions during training and then performs gradient-guided optimization at test time by iteratively noising the current solution and denoising while guiding towards lower energy solutions of some relaxed objective, achieving competitive quality–efficiency trade-offs on TSP and MIS. Fast T2T  [28] significantly speeds this up via an optimization-consistency objective, matching or surpassing multi-step diffusion solvers performance with single-step generation plus a single gradient step.
@@
Figure 2: Example of a curve (black) with three inscribed squares. Note that the inscribed squares are defined only by having all four vertices on the curve: they need not be fully contained within the curve, and they may overlap with each other.
@@
Figure 3: Inscribed square $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed (inscribed square). Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; penultimate: t=0t{=}0). For $t\neq 0$ we render only the filled mask; at t=0t{=}0 we also draw square edges and the minimum-area bounding box. The rightmost column (“snapped”) shows the rigidly snapped version of the t=0t{=}0 prediction on the curve, with an arrow from the original centroid to the snapped centroid.
Closer to our approach, some methods solve constrained problems in pixel-space representation. [20] train an unconditional diffusion model on pixel-space representations of TSP instances. They then solve new instances with stochastic optimization using a differential renderer, utilizing the prior of the learned model. Differently from us, they optimize a parametric representation of the solution to a given instance, while we generate solutions with DDIM sampling of a conditional model. [50] train a noise-prediction UNet in pixel space on a visual representation of Sudoku, which is another NP-hard combinatorial problem. They depart from fully-parallel image diffusion by (i) assigning individual noise levels to patches, and (ii) sampling patches in a learned order or hand-crafted order. They demonstrate that sampling order matters, and can substantially outperform a conditional DDPM baseline that operates fully in parallel.
3 Inscribed Square Problem
@@
The problem has since been resolved in several restricted settings. Some works show the conjecture holds for convex and piecewise analytic curves [11, 12, 13] and later results prove it for $C^{1}$-smooth curves and for curves of finite total curvature or generic $C^{1}$ curves [41, 4]. More recent work demonstrates validity under additional low-regularity assumptions [43], yet it remains open for the general Jordan curve case.
Figure 4: Solutions produced by our model. Each Jordan curve (black) is accompanied by predicted inscribed squares (colored).
@@
Algorithmic approaches for finding inscribed squares exist primarily for discrete or polygonal cases. For convex polygons, efficient procedures have been developed to detect or enumerate inscribed squares in subquadratic time [5, 40]. Discrete relaxations on grid polygons have likewise been studied and verified computationally for bounded cases [35]. All these methods represent the curve symbolically or combinatorially, relying on exact geometric descriptions. In contrast, our work formulates the problem directly in pixel space.
@@
We address the inscribed square problem by formulating it entirely in image space and training a conditional diffusion model to recover inscribed squares from noisy inputs. To this end, we generated a dataset of synthetic samples. Each sample consists of a smooth Jordan curve $C$ together with one to five inscribed squares. The curves were constructed procedurally such that they pass through the square vertices while avoiding self-intersections, and both curves and squares were rasterized into $128\times 128$ binary images. Each training example pairs a curve image with one inscribed square. Full details of the curve-generation process are provided in the supplementary material, under Section 7.1.
Our model follows the standard diffusion framework with a U-Net [37] backbone with self-attention layers [45], similar to those employed in text-to-image diffusion models such as Stable Diffusion [36]. The conditioning signal is the clean binary image of the curve, while the ground truth $x_{0}$ is the clean image of the square, both represented in two-dimensional pixel space. The conditioning image is concatenated as an additional channel to the noisy input $x_{t}$ at each timestep, effectively treating the curve as a non-noisy color channel. We train with 100 denoising steps using a standard noise schedule and mean-squared error objective, enabling the model to transform noisy square images into valid inscribed squares consistent with the given curve.
@@
$\mathcal{A}(S,C)\;=\;-\frac{1}{4}\sum_{p\in V(S)}\operatorname{dist}(p,C),$ (1)
@@
We then apply a small rigid transformation $(R_{\theta},t)$ to $\widehat{S}$ and select the configuration that maximizes this alignment score. In practice, we approximate this optimization with a coarse grid search over small rotations and translations (see Section 7.2 for details). This procedure nudges the predicted square so that its corners sit more closely on $C$, as visualized in the rightmost column of Fig. 3.
@@
Table 1: Evaluation Results. We report alignment $\mathcal{A}(S,C)$ and squareness $\mathcal{Q}$ under three conditions: before snapping, after snapping, and ground truth (GT).
@@
$\mathcal{Q}(S)\;=\;\frac{\operatorname{area}(S)}{w\cdot h}\;\cdot\;\exp\!\left(-\,2\,\Bigl|\tfrac{\max(w,h)}{\min(w,h)}-1\Bigr|\right).\vskip-2.0pt$ (2)
Figure 5: Example of a Steiner Minimal Tree. Left: The input terminal nodes, colored in red. Right: The Steiner Minimal Tree, where auxiliary Steiner points are colored in dark gray.
@@
4 Steiner Tree Problem
The second problem we cover in our case study is that of the Steiner Tree Problem. The Steiner tree problem asks, given a set of terminal points, to find a network of minimum total length that connects all terminals, where the construction is allowed to introduce additional points (Steiner points) to reduce length. In the Euclidean variant, which we focus on, Steiner points may be placed anywhere in the plane. The Steiner formulation is central to many applications where minimizing connection cost is critical. Some typical use cases for it include telecommunication  [47], PCB routing  [6], and in infrastructure layout (roads, pipelines)  [39, 8].
Figure 6: Steiner tree $x_{0}$ predictions across denoising steps. Each row corresponds to a different seed. Columns show selected $x_{0}$ predictions for decreasing timesteps $t$ from left to right (leftmost: t=Tt{=}T; rightmost: t=0t{=}0). Input points are overlaid in red.
@@ 4 Steiner Tree Problem
Formally, given a finite set of terminals $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{2}$, the Euclidean Steiner Tree (EST) problem asks for a straight-line embedded tree $T=(V,E)$ with $P\subseteq V$ that minimizes total Euclidean length $L(T)=\sum_{uv\in E}\|u-v\|_{2}$. Vertices in $V\setminus P$ are Steiner points. An optimal solution is called a Steiner minimal tree (SMT) and its length is denoted $L^{\star}(P)$ [18]. Figure  5 illustrates an instance of the problem and its optimal solution.
Optimal Ours Delta Optimal Ours Delta Optimal Ours Delta Optimal Ours Delta
Figure 7: Visual comparison of Steiner tree solutions. Each example is shown as a triplet of images: the optimal solution (Optimal), our model’s solution (Ours), and the difference between them (Delta). Input points are overlaid as red circles.
@@
For general $n$ (the number of initial given points), the EST problem is NP-hard already in the plane [17]. However, there is a polynomial-time approximation scheme (PTAS) for Euclidean Steiner tree in fixed dimensions: for any fixed $\varepsilon>0$ one can compute a $(1+\varepsilon)$-approximate tree in time polynomial in $n$ (for fixed $\varepsilon$ and dimension).
@@
SMTs in the plane are highly structured: (i) no two edges cross; (ii) each Steiner point has degree exactly $3$ and the three incident edges meet at $120^{\circ}$; (iii) all angles in the tree are at least $120^{\circ}$ (at terminals that have degree >1>1); (iv) the number of Steiner points is at most n2n-2; and (v) all Steiner points lie in the convex hull of $P$ [22, 3]. These properties provide strong geometric constraints that are exploited by both exact and approximate methods, as discussed next.
@@
Exact algorithms typically rely on generating locally optimal full Steiner trees and then selecting a minimum-length subset, with implementations such as GeoSteiner solving large 2D instances [49, 24]. On the approximation side, PTASes based on dissection or guillotine methods provide near-optimal guarantees in fixed dimensions [2, 32].
@@
While classical exact/approximate algorithms dominate practice, there have been several attempts to use learning-based solvers for the problem. For the Euclidean case, Wang et al. propose Deep-Steiner, which casts SMT construction as a sequential decision process: it discretizes the continuous search space, prunes candidates via KNN/MST neighborhoods, and then uses an attention-based policy trained with REINFORCE to add Steiner points iteratively [48]. For non-Euclidean variants (rectilinear SMT and graph STP), several learning-based methods have been explored, including RL and GNN-driven solvers and mixed neural–algorithmic pipelines [29, 30, 25, 1, 10, 34].
@@
We generate a synthetic dataset by sampling random points (between 10 and 20 for each instance), and finding their SMT using the GeoSteiner solver [24]. Each solution is then rasterized into a grayscale image with different values for edges, nodes and background. The full details for data generation are found in Section 7.3. We employ the same U-Net backbone as in the inscribed square experiment. The conditioning input consists of the rasterized terminal points, concatenated as an additional channel to the noisy input $x_{t}$ at each denoising step.
Table 2: Steiner Tree Evaluation Results. Comparison of our method, MST, and random solutions. Reported are valid tree rates and mean Euclidean length ratios ($\pm$ std) relative to the optimal solution across input point ranges.
[…]
@@
Optimal Ours Delta Optimal Ours Delta Optimal Ours Delta Optimal Ours Delta
Figure 8: Qualitative comparison of maximum-area polygon solutions. Each example is presented as a horizontal triplet: the optimal polygon (Optimal), the polygon generated by our model (Ours), and the difference (Delta). Red and blue regions in the difference maps indicate areas unique to the optimal and generated polygons, respectively. Even when visual discrepancies occur, the exclusive regions are typically of similar total area, resulting in solutions of comparable area. Input points are overlaid as red circles on both Optimal and Ours.
@@
5 Maximum Area Polygon Problem